OpenStudy (anonymous):

The population of organism A doubles every 3 hours, while population of organism B triples every 5 hours. Suppose the two organisms have equal populations at the start. When will the population of organism A be 10,000 more than that of organism B? A little help pls :) I already made an equation of this, sort of, but i just dont know how to simplify or something, it's confusing. Or maybe i did it wrong, so pls anyone? HELP :)

OpenStudy (anonymous):

mateaus ur back:)

OpenStudy (anonymous):

yey someone's replying! haha :)

OpenStudy (anonymous):

3y=2x for every 3 hours (y) = A doubles = 2x. 3(1)=2(1) --> 3=2 , 3(2)=2(2) --> 6=4...

OpenStudy (anonymous):

5y=3x for the second one and then find when A will be 10k more than B

OpenStudy (anonymous):

@Mateaus there's a formula for this right?

OpenStudy (dumbcow):

hmm no this involves exponential growth \[\large 2^{t/3} = 3^{t/5} +10,000\]

OpenStudy (anonymous):

@dumbcow i got it like that but the only difference is mine's right side is: 3^t/5 x 10, 000. haha i think im wrong though

OpenStudy (anonymous):

i equated it from the initial population.

OpenStudy (dumbcow):

wait do we know what the initial population is ? that will affect the answer

OpenStudy (dumbcow):

also is there a typo...is it "10,000 more" or "10,000 times more" ?

OpenStudy (anonymous):

\[from P = Po (2)^{t/3}\] \[i did \it like this: PO = \frac{ P }{ 2^{t/3} }\] and did the same for organism B then equated both equations :)

OpenStudy (anonymous):

they have the same initial pop :)

OpenStudy (anonymous):

@dumbcow they have the same initial population :)

OpenStudy (dumbcow):

right but what is that initial population?

OpenStudy (anonymous):

@dumbcow sorry for the typo!! haha, it's "start"

OpenStudy (dumbcow):

?? no ... is it "10,000 more" or "10,000 times more" ?

OpenStudy (anonymous):

10,000 more

OpenStudy (anonymous):

ohhhhh so im wrong bcoz i multiplied it!

OpenStudy (dumbcow):

right and it makes it lot harder to solve

OpenStudy (anonymous):

what's running in my mind is "10K more" woo haha

OpenStudy (anonymous):

uhmmm right :)

OpenStudy (anonymous):

i will try again

OpenStudy (dumbcow):

you cant solve it algabraically, i suggest using newtons method for finding zeros or wolframalpha.com

OpenStudy (anonymous):

or a graphic calculator

OpenStudy (dumbcow):

do you know answer by chance, that may clear up this confusion on setting up the equation

OpenStudy (dumbcow):

also if you are adding the 10,000 you need to know what the initial population is

OpenStudy (anonymous):

@dumbcow im using logarithm or natural log :)

OpenStudy (anonymous):

our teacher is allowing us to use scientific calculator

OpenStudy (dumbcow):

wont work if adding the 10,000....now im thinking it should be multiplied :{ can you check the wording of question in book again

OpenStudy (anonymous):

okay :) wait.

OpenStudy (anonymous):

@dumbcow double checked it. that's really the question..

OpenStudy (anonymous):

and our teacher didn't mention to us what the answer is :(

OpenStudy (dumbcow):

ok then they wrote it wrong....it should say "10,000 times more" \[2^{t/3} = 10,000*3^{t/5}\] this you can solve by using logs...the initial pop cancels

OpenStudy (anonymous):

ohh that's what i did at first because what's running in my mind was "10K more" Hahaha

OpenStudy (anonymous):

as i said, (below the problem i submitted) i came up with an equation in which i dont know how or it's hard to simplify, and that was THAT equation. :)

OpenStudy (anonymous):

my problem is about the two bases of the both with exponents

OpenStudy (anonymous):

i dont rly have the idea, i dont know, my brain is empty hahaha

OpenStudy (dumbcow):

haha ok we've come full circle...anyway \[\frac{t}{3} \ln 2 = \frac{t}{5} \ln 3 + \ln 10,000\] did you get this far

OpenStudy (anonymous):

i didnt hahaha, i was stuck the one said lol

OpenStudy (anonymous):

oh!! natural log.

OpenStudy (dumbcow):

right and 2 rules of logs are \[\ln (a*b) = \ln a + \ln b\] \[\ln (x^{n}) = n \ln x\]

OpenStudy (anonymous):

ohh i forgooot lol

OpenStudy (anonymous):

i lack knowledge in natural log :( rly.

OpenStudy (anonymous):

@dumbcow uhm, can you simplify your equation for me plss :D

OpenStudy (dumbcow):

you will need to learn those rules before the test...almost every exponential equation is solved using logs

OpenStudy (anonymous):

yah i really need to :(

OpenStudy (dumbcow):

ok move the like terms to 1 side, then factor out variable "t" \[t(\frac{\ln 2}{3} - \frac{\ln 3}{5}) = \ln 10,000\] last step is divide \[t = \frac{\ln 10,000}{\frac{\ln 2}{3} - \frac{\ln 3}{5}}\]

OpenStudy (anonymous):

sci cal! haha

OpenStudy (anonymous):

= 813.16 (sci cal )

OpenStudy (anonymous):

@dumbcow so is it, 813.16 hours? :)

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

@dumbcow Thank you very much for your time!! :) Hope u have a nice day!! :D

OpenStudy (dumbcow):

yw , you as well good luck with your class

OpenStudy (anonymous):

thanks :)